OpenStudy (anonymous):

Solve: 1/b = (6/5b) + 1

OpenStudy (anonymous):

Times both sides by b 1 = 6/5 + b 1-6/5 = b

OpenStudy (mathstudent55):

You wrote this: \(\dfrac{1}{b} = \dfrac{6}{5}b + 1\) Is that what you meant or did you mean this? \(\dfrac{1}{b} = \dfrac{6}{5b} + 1\)

OpenStudy (anonymous):

Sorry, the second one, 6 over 5 times b, plus one.

OpenStudy (mathstudent55):

Then look at @whereismymind49 's answer bec it's correct. You just need to finish by subtracting 1 - 6/5